An object is moving in an easterly direction at 4.5 m/s when it is suddenly struck so that it moves in a direction 60 degrees south of east . This force acts on the object for 5 seconds causing it to accelerate at 0.5 m/s^2 . At what speed is the object moving when the force no longer acts on it ?


1
Expert's answer
2021-10-04T17:36:33-0400

By momentum conservation law:

"mv_1=(mat+mv_2)cos60\degree"

"v_1=(at+v_2)cos60\degree"

"v_1" is speed in an easterly direction,

a is acceleration,

t is time of acting force,

v2 is final speed


"v_2=v_1/cos60\degree-at=2\cdot4.5-0.5\cdot5=6.5" m/s